# Differentiable Grouped Feedback Delay Networks for Learning Coupled Volume Acoustics

### v1

@@ IV-C2 Colouration and Sparsity Losses
where $\lambda$’s are the weighting factors for each loss term.
@@ V Subband Processing with Parallel GFDNs
(a) Parallel subband processing – training pipeline. The dashed lines represent information from the dataset to the DiffGFDNs. The reference room transfer functions (RTFs) are first passed through a reconstructing full-octave filter bank. Then, each of the $B$ filtered subband RTFs is sent to the loss function calculator. Each subband DiffGFDN’s transfer function is passed through the same filters as the RTFs before being sent to the loss function calculator.
(b) Parallel subband processing – inferencing. The incoming signal is first passed through a filter bank and each of the $B$ subbands is processed by a trained subband DiffGFDN. Additional source-receiver information is provided to the DiffGFDNs (shown in dashed lines). The outputs of each DiffGFDN are summed to get the final reverberated output.
@@
To modulate the EDC amplitudes, we can process each frequency band with a separate DiffGFDN with source and receiver gains instead of using a single DiffGFDN with source and receiver filters. We do this because using source-receiver filters will not yield equivalence between the desired and achieved RIRs (see Sec. III-A). Moreover, learning the parameters of frequency-sampled, differentiable IIR filters using deep learning presents several challenges. Depending on the chosen filter design and activation functions, the parameter space can be either overly large or too constrained [39, 40].
In [24], the authors learned the parameters of differentiable parametric equalisers (PEQs) that consist of a cascade of biquad filters parametrised as SVFs. We tried a similar parameterisation, and fixed the centre frequencies of the PEQs at octave bands, and attempted to learn the resonance and command gains of each biquad with the MLP. This resulted in large EDC errors in lower frequencies. Therefore, we propose an alternate architecture where each frequency band of a signal is processed with a different GFDN, all of whose parameters are frequency-independent.
@@
The training pipeline, shown in Fig. 2a, uses a perfect-reconstruction FIR filter bank [41] to split each measured RIR into $B$ frequency bands. The outputs of the individual GFDNs are passed through the corresponding filters before computing the loss.
At inference time, the input signal is similarly decomposed into $B$ subbands using the same filter bank. Each subband is processed by its corresponding DiffGFDN in parallel, and the outputs are summed to produce the reverberated signal. This inference pipeline is illustrated in Fig. 2b. The final transfer function of the system, $\hat{H}(\mathbf{x},z)$, is given by,
@@
where $G_{b}(z)$ is the transfer function of the $b^{\text{th}}$ subband filter, and $\bm{\Gamma}_{k,b}=\text{diag}\left(\bm{\gamma}_{k,b}\right)$ is the diagonal matrix of absorption gains, derived from the common decay times, $T_{60_{k,b}}$, using (2).
Using (8), (12) and (13), we can write the impulse response of the parallel subband GFDNs as,
@@ V-A Computational Complexity
In comparison, the CS model requires a total of $M_{b}\times B$ modes per group from (6), or a total of $G\times M_{b}\times B$ modes. Each mode can be reproduced by a biquad filter [43], which requires $9$ FLOPS. The total number of FLOPS required by the static modal reverberators is $9\times G\times M$, where $M=M_{b}\times B$. Similarly, each group requires an equaliser with $B$ cascaded biquads, so the total number of operations required by the EQs are, $9\times G\times B$. Therefore, the total number of FLOPS required by the proposed CS renderer in [16] is $9G\times(M+B)$. Karjaleinen [44] studied the number of modes, uniformly distributed over the Bark scale, required for perceptually diffuse reverberation, and used $5,000$ as a reference. For $G=2,M=2,500$, the total number of modes is $5,000$ and the number of FLOPS is $45,144$ which exceeds the FLOPS required by the proposed GFDN architecture by an order of magnitude.
@@ VI Evaluation
Each GFDN in the parallel architecture has $N=12$ delay lines grouped into $G=3$ groups ($N^{\prime}=4$), with co-prime lengths resulting in time delays between 205020–50 ms for high modal density. We also experimented with $N=8$ delay lines per group, which yielded smoother reverberation at the cost of higher computational load. We report results with $N=4$ delay lines, as it is a more efficient structure that converged to similar loss values after optimisation.
@@
With a fixed source, unit-magnitude source gains are used, and only the receiver gains are learned by the MLP. The source gain is kept fixed. The learnable matrices $\mathbf{W}_{k}$ (for the block-diagonal feedback matrix) are initialized with values from $\mathcal{U}[-\frac{1}{\sqrt{N}},\frac{1}{\sqrt{N}}]$, and input/output gains from $\mathcal{U}[-\frac{1}{N},\frac{1}{N}]$. The encoder uses spatial frequencies $f_{\textrm{min}}=1$ m-1, $f_{\textrm{max}}=32$ m-1, and $N_{f}=20$ frequency components. Inputs are expanded to $B_{b}\times 6N_{f}$ for batch size $B_{b}=32$. MLPs use ReLU activations, layer normalization [45], and He initialization [46].
@@
For subband training, RIRs were decomposed into octave bands using a $4096$-tap amplitude-preserving FIR filter bank [41] [^2] Implemented in the Pyfar library: https://pyfar.org. One DiffGFDN was trained per band, with scalar receiver gains learned by an MLP. Absorption gains were fixed based on the common decay times. We trained for $15$ epochs using the loss in (20), with an EDC loss weighting of $\lambda_{\text{EDC}}=10$. All other weights were set to $1$. MLP size varied by frequency band: $1$ hidden layer with $16$ neurons for $63$ Hz and $125$ Hz bands; $5$ hidden layers with $16$ neurons for $250$ Hz and $500$ Hz bands; and $3$ hidden layers with $128$ neurons for higher bands. These were found by searching over a grid of hyper-parameters. Input and output gains $\bm{b}$ and $\bm{c}$ were normalized at every step to ensure each group’s impulse response has unit energy. The code is available at https://github.com/orchidas/DiffGFDN/ and sound examples are available at https://ccrma.stanford.edu/~orchi/FDN/GFDN/DiffGFDN/.
VI-A Spectral Colouration and Echo Density
(a) Pre-optimisation magnitude spectrum of a group in the GFDN
(b) Post-optimisation magnitude spectrum of a group in the GFDN
@@ VI-A Spectral Colouration and Echo Density
The magnitude responses, $|\hat{H}_{k}(e^{j\omega})|^{2}$, of a single group are offset by $20$ dB and shown in Fig. 3 in yellowish-green dotted lines. Fig. 3a shows the magnitude response before training the parameters of the DiffGFDN, and Fig. 3b shows the magnitude response post-training. Other colours on the plots show the magnitude response of each subband GFDN after filtering with the respective subband filter, i.e., $|\hat{H}_{k,b}(e^{j\omega})G_{b}(e^{j\omega})|^{2}$.
@@
VI-B Comparison with Common Slopes Model
(a) $f_{c}=63$Hz, RMSE = $3.7$ dB
(b) $f_{c}=125$Hz, RMSE = $0.7$ dB
(c) $f_{c}=250$Hz, RMSE = $0.4$ dB
(d) $f_{c}=500$Hz, RMSE = $0.4$ dB
(e) $f_{c}=1$kHz, RMSE = $0.3$ dB
(f) $f_{c}=2$kHz, RMSE = $0.3$ dB
(g) $f_{c}=4$kHz, RMSE = $0.6$ dB
(h) $f_{c}=8$kHz, RMSE = $0.9$ dB
@@ VI-B Comparison with Common Slopes Model
(a) $f_{c}=63$Hz, RMSE = $5.0$ dB
(b) $f_{c}=125$Hz, RMSE = $1.4$ dB
(c) $f_{c}=250$Hz, RMSE = $1.2$ dB
(d) $f_{c}=500$Hz, RMSE = $1.5$ dB
(e) $f_{c}=1$kHz, RMSE = $0.8$ dB
(f) $f_{c}=2$kHz, RMSE = $1.2$ dB
(g) $f_{c}=4$kHz, RMSE = $1.7$ dB
(h) $f_{c}=8$kHz, RMSE = $1.4$ dB
@@
VI-B1 Octave Band EDC Losses
To analyse the results on the entire grid of positions, we plotted the error in EDC matching between the reference RIRs and the DiffGFDN’s RIRs for each position and each frequency-band. A total of $80\%$ of the receiver positions were randomly chosen to be in the training set, whereas the remaining $20\%$ constituted the test set. The results are reported for all available receiver positions in the dataset. This band-wise EDC error is calculated as
@@ VI-B1 Octave Band EDC Losses
Figs. 5 and 6 show the mean absolute EDC error, $\epsilon_{\text{EDC}_{b}}(\mathbf{x})$, as a function of receiver position according to (33) for the CS model and the proposed subband DiffGFDN architecture in octave bands. The common decay times in the CS model are the same as those used to derive the absorption gains in the DiffGFDNs. The position-dependent CS amplitudes are derived using non-linear least squares by minimising the difference between the EDCs obtained from the reference RIRs and CS predicted EDCs. Note that the CS model requires measured RIRs at all positions in the dataset, whereas the DiffGFDNs were trained on RIRs from a subset of positions (training set), and interpolated the RIRs at other positions (test set).
Although the subband DiffGFDN yields consistently low EDC errors below $1.8$ dB in Fig. 6, the CS model gives a better root-mean-squared-error (RMSE) in all frequency-bands, in Fig. 5. Both methods perform poorly in the $63$-Hz band, where dominant room modes need to be modelled individually.
VI-B2 EDR Losses
(a) Subband DiffGFDN, RMSE = $3.2$ dB
(b) CS model, RMSE = $4.8$ dB
The final RIRs from the subband DiffGFDNs are synthesised by passing an impulse through the trained subband DiffGFDNs in parallel, and filtering the output of each DiffGFDN with the appropriate subband filter, before summing all the frequency bands. The CS model’s RIRs are synthesised by shaping white noise in octave bands. The slope of the RIRs in each octave band is determined by the common decay times, and the initial level of the EDC in each octave-band for each RIR is determined by the position-dependent amplitudes.
@@ VI-C Spatial Sampling Requirements
TABLE I: RMSE in octave-band EDC fits over receiver positions in the test set as a function of different number of receivers used in the training set. The lowest errors in each frequency band are written in bold text.
@@ 63 Hz 125 Hz 250 Hz 500 Hz 1 kHz 2 kHz 4 kHz 8 kHz
30%  (252) $2.7223$ $1.2760$ $1.1188$ $1.4606$ $0.66$ $0.8046$ $1.6928$ $2.7421$
40%  (336) $2.3889$ $1.4103$ $1.0203$ $1.0266$ $0.9579$ $1.3006$ $1.3449$ $2.6807$
50%  (419) $2.6911$ $1.4754$ $1.2626$ $1.1999$ $0.5872$ $2.2138$ $1.5457$ $2.6379$
60%  (503) $2.484$ 1.16 $1.0593$ 1.023 591.023\,59 $0.5179$ $0.4328$ $0.49$ $2.6019$
70%  (587) $2.4674$ $1.2563$ 0.94 0.93 $0.4703$ $0.509$ 0.39 2.36
80%  (671) 2.20 $1.16$ $1.1379$ $0.9027$ $0.5009$ $0.4096$ $0.4331$ $2.6539$
90%  (755) $2.8416$ $1.2026$ $1.364$ $1.2359$ 0.40 0.30 $0.4669$ $2.7505$
The effect of spatial sampling on the EDC RMSE is shown in Table I. Different numbers and random distributions of receivers positions are used for training. For example, a train-test split of $40\%$ would mean $40\%$ of the receivers positions were in the training set, and $60\%$ were in the test set. The training is run for $20$ epochs on an NVIDIA V100 GPU. The errors are calculated for receiver positions in the test sets. In general, EDC errors decrease when more RIRs are available for training, although not significantly. The model is able to learn the amplitude mappings even when only $250$ receiver positions are used for training, with a mean EDC error of $1.6$ dB. There is a noticeable improvement in the mid- and high-frequency bands EDC fits when the number of receivers in the training set is increased from from $50\%$ to $60\%$. For all frequency bands except $1$ kHz and $2$ kHz, increasing the number of receivers in the training set from $80\%$ to $90\%$ results in decreased performance, likely due to overfitting.
VII Conclusion

### v2

@@ IV-C2 Colouration and Sparsity Losses
where $\lambda$’s are the weighting factors for the corresponding loss terms.
@@ V Subband Processing with Parallel GFDNs
((a))
((b))
@@
To modulate the EDC amplitudes, we can process each frequency band with a separate DiffGFDN with source and receiver gains instead of using a single DiffGFDN with source and receiver filters. We do this because using source-receiver filters will not yield equivalence between the desired and achieved RIRs (see Sec. III-A). Moreover, learning the parameters of frequency-sampled, differentiable IIR filters through deep learning poses several challenges. To avoid time-aliasing, longer responses must be evaluated, which can increase training time [24]. The estimated coefficients must be constrained to ensure stability [35, 34], which can increase the roughness of the loss landscape, requiring smoothing techniques and fine tuning. In addition, higher-order filters require deeper MLPs to learn their parameters [9]. Therefore, we prefer to process each frequency band separately to simplify the loss landscape and reduce overall complexity.
In [35], the authors learned the parameters of differentiable parametric equalisers (PEQs) that consist of a cascade of biquad filters parametrised as SVFs. We tried a similar parameterisation, and fixed the centre frequencies of the PEQs at octave bands, and attempted to learn the resonance and command gains of each biquad with the MLP. This resulted in large EDC errors in lower frequencies (demonstrated in Sec. VI-A). Therefore, we propose an alternate architecture where each frequency band of a signal is processed with a different GFDN, all of whose parameters are frequency-independent.
@@
The training pipeline, shown in Fig. 2(a), uses a perfect-reconstruction FIR filter bank [4] to split each measured RIR into $B$ frequency bands. The outputs of the individual GFDNs are passed through the corresponding filters before computing the loss.
At inference time, the input signal is similarly decomposed into $B$ subbands using the same filter bank. Each subband is processed by its corresponding DiffGFDN in parallel, and the outputs are summed to produce the reverberated signal. This inference pipeline is illustrated in Fig. 2(b). The final transfer function of the system, $\hat{H}(\mathbf{x},z)$, is given by,
@@
where $G_{b}(z)$ is the transfer function of the $b^{\text{th}}$ subband filter, and $\bm{\Gamma}_{k,b}=\text{diag}\left(\bm{\gamma}_{k,b}\right)$ is the diagonal matrix of absorption gains, derived from the common decay times, $T_{60_{k,b}}$, using (2). Using (8), (12) and (13), we can write the impulse response of the parallel subband GFDNs as,
@@ V-A Computational Complexity
In comparison, the CS model requires a total of $M_{b}\times B$ modes per group from (6), or a total of $G\times M_{b}\times B$ modes. Each mode can be reproduced by a biquad filter [1], which requires $9$ FLOPS. The total number of FLOPS required by the static modal reverberators is $9\times G\times M$, where $M=M_{b}\times B$. Similarly, each group requires an equaliser with $B$ cascaded biquads, so the total number of operations required by the EQs are, $9\times G\times B$. Therefore, the total number of FLOPS required by the proposed CS renderer in [21] is $9G\times(M+B)$. Karjalainen and Järveläinen [31] studied the number of modes, uniformly distributed over the Bark scale, required for perceptually diffuse reverberation, and used $5,000$ as a reference. For $G=2,M=2,500$, the total number of modes is $5,000$ and the number of FLOPS is $45,144$ which exceeds the FLOPS required by the proposed GFDN architecture by an order of magnitude.
@@ VI Evaluation
Each GFDN in the parallel architecture has $N=12$ delay lines grouped into $G=3$ groups ($N^{\prime}=4$), with co-prime lengths resulting in time delays between 205020–50 ms for high modal density. We also experimented with $N=8$ delay lines per group, which yielded smoother reverberation at the cost of higher computational load. We report results with $N=4$ delay lines per group, as it is a more efficient structure that converged to similar loss values after optimisation.
@@
With a fixed source, unit-magnitude source gains are used, and only the receiver gains are learned by the MLP. The source gain is kept fixed. The learnable matrices $\mathbf{W}_{k}$ (for the block-diagonal feedback matrix) are initialised with values from $\mathcal{U}[-\frac{1}{\sqrt{N}},\frac{1}{\sqrt{N}}]$, and input/output gains from $\mathcal{U}[-\frac{1}{N},\frac{1}{N}]$. The encoder uses spatial frequencies $f_{\textrm{min}}=1$ m-1, $f_{\textrm{max}}=32$ m-1, and $N_{f}=20$ frequency components. Inputs are expanded to $B_{b}\times 6N_{f}$ for batch size $B_{b}=32$. MLPs use ReLU activations, layer normalization [6], and He initialization [25].
@@
For subband training, RIRs were decomposed into octave bands using a $4096$-tap amplitude-preserving FIR filter bank [4] [^2] Implemented in the Pyfar library: https://pyfar.org. One DiffGFDN was trained per band, with scalar receiver gains learned by an MLP. Absorption gains were fixed based on the common decay times. We trained for $15$ epochs using the loss in (20), with an EDC loss weighting of $\lambda_{\text{EDC}}=10$ and sparsity loss weighting of $\lambda_{\text{sparsity}}=2$. All other weights were set to $1$. MLP size varied by frequency band: $1$ hidden layer with $16$ neurons for $63$ Hz and $125$ Hz bands; $5$ hidden layers with $16$ neurons for $250$ Hz and $500$ Hz bands; and $3$ hidden layers with $128$ neurons for higher bands. These were found by searching over a grid of hyper-parameters. Input and output gains $\bm{b}$ and $\bm{c}$ were normalised at every step to ensure each group’s impulse response has unit energy. [^3] The code is available at https://github.com/orchidas/DiffGFDN/ and sound examples are available at https://ccrma.stanford.edu/~orchi/FDN/GFDN/DiffGFDN/.
VI-A Ablation Studies
TABLE I: RMSE in octave-band EDC fits over receiver positions in the hold-out test set for different configurations tested in the ablation study. 8080% of the remaining dataset was used for training in all cases.
Configuration Mean subband DiffGFDN EDC error (dB) over receiver positions in test set
63 Hz 125 Hz 250 Hz 500 Hz 1 kHz 2 kHz 4 kHz 8 kHz
Full-band $8.9$ $4.40$ $4.13$ $5.13$ $4.81$ $1.88$ $2.09$ $3.11$
Colourless Prototype $8.44$ $3.97$ $4.09$ $3.47$ $2.48$ $2.94$ $3.36$ $1.67$
Coupled $4.29$ $1.5$ $1.88$ $1.51$ $1.15$ $1.61$ $2.82$ $1.41$
Decoupled (Proposed) $4.13$ $1.54$ $1.49$ $1.50$ $1.62$ $1.87$ $2.09$ $1.46$
To rigorously evaluate the architectural choices underlying the proposed model, we conducted a series of controlled ablation studies, each designed to isolate the effect of a specific modelling assumption. The number of delay lines was kept fixed across all configurations, as this parameter has minimal influence on the EDC mismatch and primarily affects perceptual attributes such as colouration and echo density. $10\%$ of the receiver positions in the dataset was used to create a hold-out test set. The remaining receiver positions in the dataset were split into $80\%-20\%$ for training and validation. The EDC errors are reported on the test set.
  1. 1.
    Full-band: The first experiment evaluates the viability of a unified broadband architecture. In this configuration, the MLP directly predicts the command gains and resonant frequencies of the biquad filters that implement the frequency-dependent receiver gains $g_{o,k}(z)$. Thus, only one broadband GFDN is employed, and its parameters, including the feedback matrix and input–output gains, are learned jointly with the MLP weights and biases. To give the model more expressive power, a single MLP with $10$ hidden layers and $64$ neurons per layer is used. This ablation tests whether a single, more expressive model can replace the subband architecture without sacrificing accuracy.
  2. 2.
    Colourless Prototype: The second experiment examines whether subband-specific structural parameters are necessary. Here, we fixed the set of delay-line lengths across all subband GFDNs and applied colourless optimisation [12] to obtain a single shared set of feedback matrix, $\bm{A}$, and input-output gains, $\bm{b},\bm{c}$. All subband GFDNs therefore possess identical colouration and echo density, and differ only through their receiver gains predicted by the MLP. An advantage of this configuration is that, after training, the MLP-predicted subband receiver gains can be consolidated into a single GEQ operating on octave-band cut-off frequencies, enabling inference with a single broadband GFDN. This study evaluates whether the benefits of subband decomposition arise from structural diversity across bands or merely from band-specific gain modulation.
  3. 3.
    Coupled: The third experiment investigates the role of inter-group decoupling in the feedback structure. In this variant, we relaxed the block-diagonal constraint on the feedback matrix by allowing non-zero off-diagonal entries, yielding a fully learnable dense feedback matrix. This modification introduces inter-group coupling in the feedback loop and breaks the correspondence with the CS formulation. This ablation assesses whether explicit inter-group coupling enhances modelling accuracy or instead degrades stability and interpretability relative to the proposed decoupled structure.
Table I summarises the subband EDC mismatch errors obtained on the test set for all configurations and compares them with the performance of the final architecture adopted in this work, referred to as Decoupled, which clearly outperforms Colourless Prototype and Full-band in all octave bands. Decoupled gives better or similar EDC fits compared to Coupled in all bands except $1$ kHz. For the Coupled architecture, the learnable coupling matrix becomes more diagonal after optimisation, i.e., starting from a dense matrix, the optimisation makes the final structure more closely resemble a diagonal matrix. This was verified by computing the energy ratio of the diagonal elements to its total energy pre- and post-optimisation. While allowing inter-group coupling can give marginal benefits, it breaks the correspondence with the CS model, and increases computational complexity significantly.
VI-B Evaluation of Decoupled Subband DiffGFDNs
VI-B1 Spectral Colouration and Echo Density
((a))
((b))
@@ VI-B1 Spectral Colouration and Echo Density
The magnitude responses, $|\hat{H}_{k}(e^{j\omega})|^{2}$, of a single group are offset by $20$ dB and shown in Fig. 3 in yellowish-green dotted lines. Fig. 3(a) shows the magnitude response before training the parameters of the DiffGFDN, and Fig. 3(b) shows the magnitude response post-training. Other colours on the plots show the magnitude response of each subband GFDN after filtering with the respective subband filter, i.e., $|\hat{H}_{k,b}(e^{j\omega})G_{b}(e^{j\omega})|^{2}$.
@@
VI-B2 Comparison with Common Slopes Model
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
@@ VI-B2 Comparison with Common Slopes Model
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
@@
((a))
((b))
Figure 7: Comparison of broadband EDC fits of the proposed subband GFDN architecture with the CS model and ground truth, for two receiver positions in the dataset. The multi-slope decay is evident in both plots.
((a))
((b))
  • Octave Band EDC Losses: To analyse the results on the entire grid of positions, we plotted the error in EDC matching between the reference RIRs and the DiffGFDN’s RIRs for each position and each frequency-band. A total of $80\%$ of the $838$ receiver positions were randomly chosen to be in the training set. The plots show the EDC error for all available receiver positions in the dataset. This band-wise EDC error is calculated as
@@
    Figs. 5 and 6 show the mean absolute EDC error, $\epsilon_{\text{EDC}_{b}}(\mathbf{x})$, as a function of receiver position according to (33) for the CS model and the proposed subband DiffGFDN architecture in octave bands. The common decay times in the CS model are the same as those used to derive the absorption gains in the DiffGFDNs. The position-dependent CS amplitudes are derived using non-linear least squares by minimising the difference between the EDCs obtained from the reference RIRs and CS predicted EDCs. Note that the CS model requires measured RIRs at all positions in the dataset, whereas the DiffGFDNs were trained on RIRs from a subset of positions (training set), and interpolated the RIRs at other positions (test set). Although the subband DiffGFDN yields consistently low EDC errors below $1.8$ dB in Fig. 6, the CS model gives a better root-mean-squared-error (RMSE) in all frequency-bands, in Fig. 5. Both methods perform poorly in the $63$ Hz band, where dominant room modes need to be modelled individually.
  • EDR Losses: The final RIRs from the subband DiffGFDNs are synthesised by passing an impulse through the trained subband DiffGFDNs in parallel, and filtering the output of each DiffGFDN with the appropriate subband filter, before summing all the frequency bands. The CS model’s RIRs are synthesised by shaping white noise in octave bands. The slope of the RIRs in each octave band is determined by the common decay times, and the initial level of the EDC in each octave-band for each RIR is determined by the position-dependent amplitudes.
    In Fig. 7, we show the EDC plots for the broadband RIR reconstructed from the subband GFDN outputs and the CS parameters against the ground truth data for two receiver positions in the second and third room respectively. The plots clearly demonstrate multi-slope decay, which is captured accurately by the proposed subband GFDN architecture.
@@ VI-C Spatial Sampling Requirements
TABLE II: RMSE in octave-band EDC fits over receiver positions in the hold-out test set as a function of different number of receivers used in the training set. The lowest errors in each frequency band are written in bold text.
@@ 63 Hz 125 Hz 250 Hz 500 Hz 1 kHz 2 kHz 4 kHz 8 kHz
10% (76) $5.9$ $3.7$ $3.4$ $3.5$ $2.9$ $3.9$ $4.2$ $1.7$
20% (151) $4.2$ $3.5$ $2.3$ $4.3$ $2.6$ $4.4$ $3.7$ $1.85$
30% (227) $4.2$ $2.7$ $2.5$ $4.1$ $1.6$ $4.2$ $3.8$ $1.4$
40% (302) $3.8$ $3.2$ $1.8$ $2.5$ $1.7$ $4.2$ $3.9$ $1.6$
50% (378) $4.8$ $2.2$ $1.8$ $2.2$ $1.8$ $3.5$ $3.7$ $1.6$
60% (453) $4.0$ $2.1$ $1.5$ $1.7$ $1.7$ 1.2 1.7 1.3
70% (529) $4.0$ $1.9$ $1.4$ $1.7$ $2.3$ $1.7$ $1.5$ $1.4$
80% (604) $4.3$ 1.7 1.3 $1.6$ 1.3 $3.8$ $1.8$ $1.4$
90% (680) 3.9 $2.0$ $1.6$ 1.5 $1.8$ $1.4$ $1.8$ $1.4$
The effect of spatial sampling on the EDC RMSE is shown in Table II. A fixed hold-out test set comprising $10\%$ of the data was reserved for evaluation, while the remaining $90\%$ of the receiver positions were randomly split into training sets of varying sizes. For example, a $40\%$ split indicates that $40\%$ of the remaining data was used for training and the rest for validation. Training was run for $20$ epochs on an NVIDIA V100 GPU. All reported errors are computed over the fixed $10\%$ test set. Errors typically decrease with more training data, as expected. Low-frequency bands, particularly $63$ Hz, exhibit the highest errors and remain difficult to learn even with large amounts of data, whereas mid-frequency bands ($250$ Hz, $500$ Hz, $1$ kHz) show steady improvement as more training RIRs are provided. High-frequency bands converge earlier, although the $2$ kHz and $4$ kHz bands display non-monotonic behaviour and require at least $60\%$ of the training data before stabilising. When using $80\%$ of the data for training, the error for the $2$ kHz band fails to converge within $20$ epochs. Across most octave bands, a noticeable reduction in error variability occurs around the $60\%$ mark, after which gains become more incremental.
VII Discussion
Ablation studies showed that the decoupled subband DiffGFDN architecture provides the best trade-off between accuracy and model complexity. Analysis of the magnitude response and NED curves further confirmed that jointly learning the feedback matrix and the input–output gains reduces timbral colouration and increases echo density. Compared to the state-of-the-art CS model, the proposed architecture yields slightly higher EDC errors but achieves lower EDR errors, while also enabling interpolation to unseen receiver positions.
We examined the effect of spatial sampling density on octave-band EDC errors using a fixed hold-out test set. As expected, using more RIR samples for training reduced the error which stabilised to values less than $3$ dB in all except the $63$ Hz band when using $60\%$ or more of the data for training. The only anomaly is the error for the $2$ kHz band when trained using $80\%$ of the data. Determining the spatial sampling requirements for perceptually robust late-reverberation modelling—including the number and spatial arrangement of necessary RIRs—remains an open research question.
A natural next step is to evaluate DiffGFDN on real measured rooms. However, the scarcity of high-quality multichannel RIR datasets poses practical challenges. Real measurements often contain noise and artefacts that are absent in simulated datasets. To operate reliably on such data, DiffGFDN would need to explicitly model the noise floor and potentially incorporate denoising into the training process. Addressing these issues and collecting suitable measurement datasets will be essential for validating the method in real-world acoustic environments.
The subband DiffGFDN architecture is well-suited for rendering late reverberation in coupled spaces for AR applications when dense RIR measurements are available. Unlike storing full RIR datasets in the cloud, the learned DiffGFDN parameters can be stored directly on-device. Our recent extension of the method to directional late reverberation using spherical microphone array measurements enables binaural rendering [18]. Rendering is orders of magnitude more efficient than time–space–varying convolution, making DiffGFDN a memory- and computation-efficient alternative to methods such as Neural Acoustical Fields [36] and Novel View Acoustic Synthesis [8]. Furthermore, the method is geometry-agnostic, relying only on RIR measurements and corresponding source–receiver positions. However, it does not yet generalise across different geometries. Few-shot generalisation has been explored for other neural RIR prediction methods [55, 28, 37] and represents a promising direction for future work with DiffGFDN.
VIII Conclusion